OpenStudy (anonymous):

Find the volume of a cube with the given side length. 10 units A. 10,000 cubic units B. 1000 cubic units C. 100 cubic units D. 30 cubic units

OpenStudy (anonymous):

is it 30?

OpenStudy (alexandervonhumboldt2):

volume=side^3

OpenStudy (alexandervonhumboldt2):

volume=10^3=10*10*10=?

OpenStudy (anonymous):

30?